# Can the 3 neutrino masses really be found using SN 1987A data?

### v3

@@ IV.4 Alternate explanations for the LSD neutrinos?
  • The only other explanation known to this author for the LSD burst besides tachyons is the Frogatt-Nielsen proposal of “dark matter balls” inside the the stellar core (Froggatt2015, ).
@@ IV.9 Consistency with direct mass experiments?
IV.12 Finding more evidence in oscillation data?
Neutrino oscillation experiments of the dissapearing type can also be used to find evidence of the three neutrino masses from the SN 1987A data using a similar approach. The key quantity fitted to the data in such an experiment is the survival probability $P_{j}$ of a neutrino of a given flavor. One usually looks to see if there is any significant amplitude $\alpha_{j}$ for the probability anywhere on a grid of values over sterile neutrino mass and amplitude of the oscillation for a single sterile neutrino.
To find how much amplitude oscillation there is in the data for neutrinos of mass $m_{1}=2.70eV/c^{2}$ and simultaneously for neutrinos of mass $m_{2}=22.4eV/c^{2}$ one can create a grid of $\alpha_{1}$ and $\alpha_{2}$ values with the summed probability $P_{1}+P_{2}$ calculated at each grid point, and see if there is a small region of the grid where the summed probabilities is very high. If so, that would constitute strong support for the idea that the $m^{2}$ masses of electron neutrinos consist of the values deduced from the SN 1987A data. The most promising experiments to look at might be the four identified in ref. (Serebrov2024, ) which already have shown evidence of the $2.70eV/c^{2}$ mass – do they now show evidence of a $22.4eV/c^{2}$ neutrino as well? Of course, only some experiments having a very short baseline will be sensitive to a mass this high. It would also be interesting to look at some formerly negative results such as one from the NOvA Collaboration to see if they see anything. (Acero2025, )
@@ V Summary and the ultimate test
This paper has made six claims that many. readers may understandably consider very difficult to believe – see Table 1. Some of the claims will likely not be accepted as true without the strongest corroborating evidence. After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data with little uncertainty in their values, that might be enough. This might also be the case if dissapearance experiments were to find such evidence doing the analysis mentioned in the previous section.
1. No valid basis exists to ignore the LSD 5-hr early burst
2. No a priori upper limits exist on the three neutrino $m^{2}_{\nu,j}$
@@ 4. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
6. Existing data likely can prove or reject all the above

### v4

@@ IV.4 Alternate explanations for the LSD neutrinos?
  • The only other explanation known to this author for the LSD burst besides tachyons is the Froggatt-Nielsen proposal of “dark matter balls” inside the the stellar core (Froggatt2015, ).
@@ IV.9 Consistency with direct mass experiments?
IV.12 Multiple sterile neutrinos in oscillation data?
Anatolii Serebrov and his Neutrino 4 team, who first detected reactor antineutrino oscillations in the 6-12 meter range, has advocated the need for two more sterile neutrinos in addition to $m_{1}^{2}=2.70^{2}eV^{2}/c^{4}$ based on theoretical grounds (Serebrov2024, ) Assuming a second sterile neutrino of mass $m_{2}^{2}=22.4eV^{2}/c^{2},$ the oscillations would presumably show up in a range roughly 80 times shorter. If such oscillations are detectable one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find thelikelihood probability of the fit to the data for that pair of mass values. One could then see where on the plane that likelihood maximizes. If there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},22.4^{2})eV^{2}/c^{2},$ that would constitute strong support for the idea that the $m^{2}$ masses of electron neutrinos consist of the values deduced from the SN 1987A data. The most promising experiments to look at might be the four identified by Serebrov in ref. (Serebrov2024, ). These already have shown evidence of the $2.70eV/c^{2}$ mass, and the question is do they now also show evidence of a $22.4eV/c^{2}$ neutrino. It would seem highly unlikely that a signature of any oscillation due to the $m_{3}^{2}=-200,000eV^{2}/c^{4}$ neutrino could be found in a similar manner, but it should be detectable in KATRIN as already discussed.
@@ V Summary and the ultimate test
This paper has made six claims that many. readers may understandably consider very difficult to believe – see Table 1. Some of the claims will likely not be accepted as true without the strongest corroborating evidence. After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data with little uncertainty in their values, that might be enough. This might also be the case if oscillation experiments were to find such evidence doing the analysis mentioned in the previous section.
1. The LSD 5-hr early burst was connected to SN 1987A
2. Upper limits on the 3 neutrino $m^{2}_{\nu,j}$ can be avoided
@@ 4. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
6. Existing data likely can confirm or reject all the above